### v4

# Fast Robot Adaptation via Hand Path Retrieval

Abstract
We hand the community HAND, a simple and time-efficient method for teaching robots new manipulation tasks through human hand demonstrations. Instead of relying on task-specific robot demonstrations collected via teleoperation, HAND uses easy-to-provide hand demonstrations to retrieve relevant behaviors from task-agnostic robot play data. Using a visual tracking pipeline, HAND extracts the motion of the human hand from the hand demonstration and retrieves robot sub-trajectories in two stages: first filtering by visual similarity, then retrieving trajectories with similar behaviors to the hand. Fine-tuning a policy on the retrieved data enables real-time learning of tasks in under four minutes, without requiring calibrated cameras or detailed hand pose estimation. Experiments also show that HAND outperforms retrieval baselines by over $2\times$ in average task success rates on real robots. Videos can be found at our project website: https://liralab.usc.edu/handretrieval/.
Equal Contribution, Equal Advising
I Introduction
For robots to operate seamlessly in human-centric settings, they should be able to rapidly learn new tasks with minimal human input. Achieving this goal requires robot learning algorithms that (1) scale across many tasks and (2) adapt quickly to new ones.
Imitation learning has produced capable multi-task robot policies [1, 2, 3, 4, 5], but scaling is hindered by its reliance on vast amounts of expert-collected, task-specific teleoperation data [6]. In contrast, task-agnostic play data is far easier to collect, without requiring constant environment resets or task-specific labeling [7, 8, 9]. The challenge is in making such unstructured data usable for teaching robots new tasks quickly.
Therefore, we propose HAND, a simple and time-efficient approach that adapts pre-trained play policies to specific tasks using just one human hand demonstration (see LABEL:fig:teaser). Unlike prior retrieval methods [10, 11, 12, 13, 14, 15] that require robot demonstrations of the target task, HAND extracts coarse, 2D relative hand motion paths from the provided human hand demonstration to guide retrieval. Thus, our approach enables even non-experts to teach robots without teleoperation.
HAND enables both scalability and speed. Towards scalability, HAND avoids the need for calibrated depth cameras [16, 17], specialized eye-in-hand setups [18], or detailed hand-pose estimation [18, 19]. Instead, it first labels a robot play dataset with 2D gripper positions relative to the RGB camera frame by tracking the gripper using a visual point-tracking model [20, 21]. When a human hand demonstration is provided, HAND tracks the hand trajectory with the same simple pipeline. The hand positions are then converted into 2D relative sub-trajectories, capturing motion independent of the starting point [22]. After an initial filtering step that removes unrelated behaviors using a visual foundation model [23], HAND retrieves matching sub-trajectories from the play dataset based on the 2D relative hand path. Finally, towards speed, a policy pre-trained on the play dataset is LoRA-fine-tuned on the retrieved sub-trajectories, encouraging the policy to specialize in the demonstrated task. Because HAND retrieves primarily based on hand motion, it is robust to irrelevant visual features such as background clutter and lighting changes compared to purely visual retrieval methods.
[…]
II Related Works
Robot Data Retrieval. Prior work has demonstrated retrieval as an effective mechanism for extracting relevant on-robot data for training robots [10, 11, 12, 13, 14, 24, 15, 25]. For example, SAILOR [10] and Behavior Retrieval [11] pre-train variational auto-encoders (VAEs) on prior robot images and actions to learn a latent embedding. This latent embedding is used to retrieve states and actions from an offline dataset similar to ones provided in expert demonstration trajectories. However, retrieving based on learned full image encodings or even raw pixel values [14] can be noisy; Flow-Retrieval [12] instead trains a VAE to encode optical flows indicating movement of objects and the robot arm in the scene. Similar to Flow-Retrieval, our method HAND also retrieves based on robot arm movement. However, rather than training a dataset-specific VAE model that may not be robust to large visual differences, we retrieve from our offline robot data by primarily matching motions of a human hand demonstration using relative 2D paths of the robot end-effector in the prior data. This hand path retrieval helps us robustly retrieve relevant robot arm behaviors.
STRAP [13] addresses visual retrieval robustness issues of prior work by using features from DINO-v2 [23], a large pre-trained image-input foundation model for retrieval. However, STRAP, along with all aforementioned retrieval work, assumes access to expert robot demonstrations for the target tasks. HAND, on the other hand, only requires a single, easier-to-collect human hand demonstration that results in more time-efficient learning of demonstrated tasks compared to methods requiring robot teleoperation data for retrieval. Moreover, experiments demonstrate HAND actually retrieves more task-relevant trajectories and therefore attains higher success rates compared to these methods.
Learning From Human Hands. Similar to HAND, a separate line of work has proposed methods to use human hands to learn robot policies. One approach is to train models on human video datasets to predict future object flows [26, 27] or human affordances [28, 29]. These intermediate affordance and flow representations are then used to either train a policy conditioned on this representation [26] on robot data or control a heuristic policy [27, 28, 29]. Other works focus on learning directly from human hands [16, 17, 18, 30, 19]. These works generally use hand-pose detection models aided by multiple cameras or calibrated depth cameras to convert hand poses directly to robot gripper keypoints [16, 17, 19]. However, works that exclusively retrieve human data are restricted to constrained policy representations as they must match human hand poses to robot gripper poses. [18] instead use an eye-in-hand camera mounted on a human demonstrator’s forearm to train an imitation learning policy conditioned on robot eye-in-hand camera observations. Unlike these prior works, HAND only requires a single RGB camera from which the robot gripper can be seen. Also, we focus on retrieving robot play data, allowing us to train arbitrarily expressive policies without constrained policy representations [16, 17, 19] or intermediate representations [26, 27, 28, 29].
III HAND: Fast Robot Adaptation
via Hand Path Retrieval
Figure 2: HAND enables fast-adaptation to a new target task by using an easy-to-provide hand demonstration of the target task (Left). We propose a two-step retrieval procedure where we first filter the trajectories in the offline play dataset, $\mathcal{D}_{\text{play}}$, for visually similar trajectories based on features from a pretrained vision model. We use off-the-shelf, pretrained hand detection and point tracking to construct 2D paths of the motion for both the human hand and robot end-effector. We use these paths as a distance metric to retrieve relevant trajectories from the play dataset (Middle) for quickly fine-tuning a pretrained transformer policy on the target task (Right).
III-A Preliminaries and Formulation.
We assume access to a dataset of task-agnostic robot play data, $\mathcal{D}_{\text{play}}$, consisting of trajectories $\tau_{i}=\{(o_{t},a_{t})\}_{i=1}^{T}$}, where each $o_{t}$ is per-timestep observation that includes RGB images of the robot gripper and robot proprioceptive information, and $a_{t}$ is the robot action. These trajectories may span many scenes, tasks, and time horizons. We do not assume task labels (e.g., language labels), as data collection is easier to scale without labeling each sub-trajectory in a long-horizon play trajectory. [^1] Section IV demonstrates that HAND can also incorporate language labels as extra policy conditioning.
In contrast to retrieval methods that rely on robot demonstrations for each target task [10, 11, 12, 13], we assume access to easy-to-provide human hand demonstrations. For each task, a human records their hand movement without teleoperating the robot. On our real-world setup, these hand demonstrations, $\mathcal{D}_{\text{hand}}$, are on average $5\times$ faster to collect than robot teleoperation data. Moreover, producing high-quality hand demonstrations typically requires far less effort than robot teleoperation [31, 32]. Each video in $\mathcal{D}_{\text{hand}}$ consists of a sequence of RGB images $o_{1},\ldots,o_{H}$, captured from a similar viewpoint relative to the human hand as trajectories in the robot play data relative to the robot gripper. [^2] Section IV demonstrates HAND works under large camera angle shifts.
Given $\mathcal{D}_{\text{play}}$ and $\mathcal{D}_{\text{hand}}$, we aim to train a policy $\pi_{\theta}(a\mid o)$ to perform the target task demonstrated by the human in $\mathcal{D}_{\text{hand}}$. Since we do not assume task labels in $\mathcal{D}_{\text{play}}$ and we are provided no expert robot teleoperation demonstrations, we must retrieve sub-trajectories indicating how to perform the behavior demonstrated in $\mathcal{D}_{\text{hand}}$ from $\mathcal{D}_{\text{play}}$ for training $\pi$. We denote this retrieved dataset, later used for imitation learning, as $\mathcal{D}_{\text{retrieved}}$. Moreover, following our motivation in Section I, we aim for our method to be fast, so that non-expert end-users can easily train the robot for many downstream tasks.
Thus, the key challenges we resolve in our method HAND are: (1) designing a representation that can unify the behaviors in robot sub-trajectories and human hand demonstrations (Section III-B), (2) retrieving relevant sub-trajectories based on a suitable distance metric between these representations (Section III-C), and (3) quickly training a policy that can perform various unseen target tasks with a high success rate without expert demonstrations (Section III-D). See Figure 2 for an overview.
III-B Path Distance as a Unifying Representation for Retrieval
Prior robot retrieval methods assume access to expert demonstrations from which they extract proprioceptive information (e.g., joint angles and actions) alongside visual features for retrieval [10, 11, 12, 13, 14]. However, since $\mathcal{D}_{\text{hand}}$ contains only visual data and no robot actions, retrieval based purely on appearance can be noisy—especially due to the visual domain gap between hand demonstrations in $\mathcal{D}_{\text{hand}}$ and robot demonstrations in $\mathcal{D}_{\text{play}}$ (c.f., Figure 2, left). To address these issues, we propose an embodiment-agnostic, behavior-centric retrieval metric that enables matching between $\mathcal{D}_{\text{hand}}$ and $\mathcal{D}_{\text{play}}$ based on demonstrated behaviors rather than appearance.
Using 2D Paths for Retrieval. The movement of the robot end-effector over time provides rich information about its behavior [4]. We represent behaviors in both datasets using the paths traced by the human hand or the gripper. Because we assume access only to an RGB camera from which the hand or the gripper is visible (i.e., no depth), we construct these paths in 2D relative to the camera viewpoint for both $\mathcal{D}_{\text{play}}$ and $\mathcal{D}_{\text{hand}}$. [^3] If both datasets have additional calibrated depth information, HAND can also operate on 3D paths.
Obtaining Paths from Data. To extract paths, we use CoTracker3 [21], an off-the-shelf point tracker capable of tracking 2D points across video sequences, even under occlusion. CoTracker3 only requires a single point on the gripper or hand to generate a complete trajectory. We use Molmo-7B [33], an open-source 7B image-to-point foundation model, to automatically select this point by prompting it at the midpoint of each trajectory with either “Point at the center of the hand” or “Point to the robot gripper.” Using the middle frame ensures a higher chance of visibility in case the gripper or hand is not yet in frame at the beginning or occluded at the end. [^4] Points can also be obtained heuristically, e.g., if the robot starts from the same position in each $\mathcal{D}_{\text{play}}$ traj.
Given the 2D point $(x,y)_{\text{hand}}$ or $(x,y)_{\text{play}}$ from the middle frame, we use CoTracker3 to perform bidirectional point tracking, resulting in a 2D path $p_{\text{hand}}=\{(x_{t},y_{t})_{\text{hand}}\}_{t=1}^{H}$ or $p_{\text{play}}=\{(x_{t},y_{t})_{\text{play}}\}_{t=1}^{T}$ for each trajectory. See the Gripper/Hand Tracking block of Figure 2 for a visualization of this pipeline. Next, we describe how we use 2D paths to retrieve sub-trajectories from $\mathcal{D}_{\text{play}}$.
III-C Retrieving Relevant Sub-Trajectories using Path Distance
Background. For identifying relevant sub-trajectories in $\mathcal{D}_{\text{play}}$, we use Subsequence Dynamic Time Warping (S-DTW) [34], an algorithm for aligning a shorter sequence to a portion of a longer reference sequence prior work has demonstrated effective for sub-trajectory retrieval [13]. Given a query sequence $Q=\{q_{1},q_{2},\dots,q_{H}\}$ and a longer reference sequence $R=\{r_{1},r_{2},\dots,r_{T}\}$, where $T>H$, the goal of S-DTW is to find a contiguous subsequence of $R$ that minimizes the total cumulative distance between elements of both sequences. In HAND, the query sequences are the 2D hand demonstration paths $\{(x_{t},y_{t})_{\text{hand}}\}_{t=1}^{H}$ and the reference sequences are the 2D paths generated from long-horizon robot play data $\{(x_{t},y_{t})_{\text{play}}\}_{t=1}^{T}$.
Sub-Trajectory Preprocessing. To preprocess the datasets for S-DTW, we first segment the offline play dataset, $\mathcal{D}_{\text{play}}$, into variable-length sub-trajectories using a simple heuristic based on proprioception proposed in several prior works [35, 13]. In particular, we split the trajectories whenever the acceleration or velocity magnitude (depending on what proprioception data is available) drops below a predefined $\epsilon$ value, corresponding to when the teleoperator switches between tasks. We find that this simple heuristic can reasonably segment trajectories into atomic components resembling lower-level primitives. We also split the hand demonstrations evenly into smaller sub-trajectories based on how many subtasks the human operator determined they have completed. After sub-trajectory splitting, we have two sub-trajectory datasets, $\mathcal{T}_{\text{hand}}=\{t_{1:a}^{i},t_{a:b}^{i},\dots,t_{H_{i}-|p_{\text{hand}}^{i}|:H_{i}}^{i}\forall\,\tau^{i}_{\text{hand}}\in\mathcal{D}_{\text{hand}}\}$} and $\mathcal{T}_{\text{play}}=\{t_{1:a}^{j},t_{a:b}^{j},\dots,t_{T_{j}-|p_{\text{play}}^{j}|:T}^{j}\forall\,\tau^{j}_{\text{play}}\in\mathcal{D}_{\text{play}}\}$} where $|p_{\text{hand}}^{i}|$ and $|p_{\text{play}}^{j}|$ are the lengths of the last sub-trajectory paths of trajectories $i,j$ from $\mathcal{D}_{\text{hand}}$ and $\mathcal{D}_{\text{play}}$, respectively. Finally, each sub-trajectory is represented in relative 2D coordinates, i.e., $p_{t}=[x_{t+1}-x_{t},y_{t+1}-y_{t}]$}. Relative coordinates ensure retrieval invariance to the initial positions of the hand or gripper [22].
Visual Filtering. One issue with retrieving sub-trajectories based only on path distance is that different tasks can have similar movement patterns. For example, tasks like “pick up the mug” and “pick up the cube” can appear nearly identical in 2D path space [4]. But, the retrieved trajectories for one task may not benefit learning of the other; since we do not assume task labels in $\mathcal{D}_{\text{play}}$, a policy directly trained on “pick up the cube” retrieved sub-trajectories may still fail to pick up a mug. Therefore, before retrieving sub-trajectories with paths, we first run a visual filtering step to ensure that the sub-trajectories we retrieve will be task-relevant. We use an object-centric visual foundation model, namely DINOv2 [23], to first filter out sub-trajectories performing unrelated tasks with different objects. Specifically, we use the DINOv2 first and final frame embedding differences, representing visual object movement from the first to last frame, between human hand demonstrations and robot play data to filter $\mathcal{T}_{\text{play}}$. We find that using this simple method is sufficient to filter out most irrelevant sub-trajectories. For a given image sequence $o_{1:H}^{\text{hand}}$ from a hand sub-trajectory and image sequence $o_{1:T}^{\text{play}}$ from a robot play sub-trajectory, we define the cost as:
[…]
We take the $M$ trajectories with lowest cost as possible retrieval trajectories from $\mathcal{D}_{\text{play}}$ for each human demonstration sub-trajectory in $\mathcal{T}_{\text{hand}}$. The rest are ignored for those hand demonstrations.
Retrieving Sub-Trajectories. Finally, we then employ S-DTW to match the target sub-trajectories, $\mathcal{T}_{\text{hand}}$, to the set of visually filtered segments $\in\mathcal{T}_{\text{play}}$. Given two sub-trajectories, $t_{i}\in\mathcal{T}_{\text{play}}$ and $t_{j}\in\mathcal{T}_{\text{hand}}$, S-DTW returns the cost along with the start and end indices of the subsequence in $t_{j}$ that minimizes the path cost (see Figure 2). We select the $K$ matches from $\mathcal{D}_{\text{play}}$ with the lowest cost to construct our retrieval dataset, $\mathcal{D}_{\text{retrieved}}$.
III-D Putting it All Together: Fast-Adaptation with Parameter-Efficient Policy Fine-tuning
We aim to enable fast, data-efficient learning of the task demonstrated in $\mathcal{D}_{\text{hand}}$. To this end, we first pretrain a task-agnostic base policy $\pi_{\text{base}}$ on $\mathcal{D}_{\text{play}}$ with standard behavior cloning (BC) loss. While our approach is compatible with any policy architecture, we use action-chunked transformer policies [36] due to their suitability for low-parameter fine-tuning and strong performance in long-horizon imitation learning [37, 38, 39, 3].
Adapting to $\mathcal{D}_{\text{retrieved}}$. To rapidly adapt to a new task with minimal data, we leverage parameter-efficient fine-tuning using task-specific adapters—small trainable modules that modulate the behavior of the frozen base policy. Adapter-based methods have shown promise in few-shot imitation learning [40, 41], making them ideal for our limited retrieved dataset $\mathcal{D}_{\text{retrieved}}$. Specifically, we insert LoRA layers [42] into the transformer blocks of $\pi_{\text{base}}$. These are low-rank trainable matrices (about $0.1\%$$2\%$ of $\pi_{\text{base}}$’s parameters) inserted between the attention and feedforward layers (see Figure 2, LoRA Layers). During fine-tuning, we update only the parameters of these LoRA layers, $\theta$, using $\mathcal{D}_{\text{retrieved}}$.
Loss Re-Weighting. While our retrieval mechanism identifies sub-trajectories relevant to the target task, not all will be equally useful. Following prior work [14, 15, 25], we reweight the BC loss with an exponential term $\in(0,\infty)$ (similar to AWR [43]), where each sub-trajectory is weighted based on its S-DTW similarity to the hand demonstration. Intuitively, this upweights the loss of the most relevant examples in $\mathcal{D}_{\text{retrieved}}$ and downweights those that are less relevant. Finally, because trajectory cost scales vary depending on the task being retrieved and the features being used for S-DTW, we rescale the S-DTW costs $C_{i,\text{path}}$ to a fixed range. For each $\tau_{i}\in\mathcal{D}_{\text{retrieved}}$, its weight $e^{-C_{i,\text{path}}}$ is scaled to between $[0.01,100]$, where the normalization term comes from the sum of costs of all trajectories in $\mathcal{D}_{\text{retrieved}}$. Let the normalized weight for a trajectory be $w_{i}=\exp(-C_{i,\text{path}})$ and the behavioral cloning loss be $L_{i}(a,o)=-\log\pi_{\theta}(a\mid o)$. The total loss is then the weighted average over the dataset $\mathcal{D}$:
$\mathcal{L}_{\text{BC};\theta}=\frac{1}{|\mathcal{D}_{\text{retrieved}}|}\sum_{\tau_{i}\in\mathcal{D}_{\text{retrieved}}}w_{i}\times L_{i}(a,o).$ (2)
We summarize HAND in the pseudocode in Algorithm 1.
Algorithm 1 HAND Pseudocode
[…]
2:Train base policy $\pi_{\text{base}}$ on $\mathcal{D}_{\text{play}}$ via behavior cloning
3:Segment both $\mathcal{D}_{\text{play}}$ and $\mathcal{D}_{\text{hand}}$ into sub-trajectory datasets w/ threshold $\epsilon$: $\mathcal{T}_{\text{play}}$, $\mathcal{T}_{\text{hand}}$
4:for $\tau^{\text{hand}}\in\mathcal{T}_{\text{hand}}$ do
5:  Filter top-$M$ visually similar $\tau^{\text{play}}\in\mathcal{T}_{\text{play}}$ via DINO-based $\text{C}_{\text{visual}}$
6:  for each filtered $\tau^{\text{play}}$ do
7:   Track 2D hand paths with Molmo + CoTracker3
8:   Retrieve $K$ best-matching segments via S-DTW on relative path similarity
9:Fine-tune $\pi_{\text{base}}$ on retrieved data with adapter layers $\theta$ to obtain $\pi_{\theta}$ with $\mathcal{L}_{\text{BC};\theta}$ Equation 2
10:return $\pi_{\theta}$
IV Experiments
Our experiments demonstrate the efficacy of HAND as a robot data retrieval pipeline and evaluate its ability to quickly learn to solve downstream tasks. To this end, we organize our experiments to answer the following questions, in order:
  1. (Q1)
    How well can HAND retrieve task-relevant behaviors?
  2. (Q2)
    Does HAND support hand demonstrations from unseen scenes and is it robust to visual shifts?
  3. (Q3)
    How does HAND perform in policy learning?
  4. (Q4)
    Can HAND enable real-time adaptation?
IV-A Experimental Setup
We evaluate HAND on a real-world multi-task kitchen environment using the WidowX robot arm. Our robot environment setup is shown in Figure 3. We use an Intel Realsense D435 camera as an external camera and a Logitech C920 as an over-the-shoulder camera.
Evaluation Tasks: We first evaluate on 10 total tasks. We first evaluate on three standard tasks: Reach Green Block, Press Button, and Close Microwave. Then, we introduce three challenging long-horizon tasks: Put K-Cup in Coffee Machine, Blend Carrot, and Cook Carrot, which demand high precision and span more than 150 timesteps at a 5 Hz control frequency. In particular, Cook Carrot is composed of four shorter tasks, Slide Pot $\rightarrow$ Put Object in Pot $\rightarrow$ Put Lid on Pot $\rightarrow$ Turn Stove Knob, including non-prehensile tasks (e.g., slide pot) and taking $\sim 300$ steps to complete even for expert teleoperators. For our long-horizon tasks, we provide one hand demonstration to perform retrieval for each subtask. These tasks highlight the ability of HAND to acquire and execute complex behaviors in real time. Partial success is provided for tasks composed of multiple subtasks.
Play Dataset Collection: We collect a task-agnostic play dataset containing a total of 50k transitions, each trajectory having an average of 230 timesteps and covering multiple tasks, collected at 5hz. The full dataset required roughly four hours to collect. We place distractor objects not used in target tasks in the environment that teleoperators interact with during play data collection to ensure the play data does not mirror evaluated tasks. The dataset is split into two, with about 1 hour corresponding to the scene for Cook Carrot. To evaluate language-conditioned methods, we manually annotate the Cook Carrot scene of the dataset with language, which takes an additional 87 minutes. During data collection and evaluation, movable task objects are randomized in a 5” x 7” region within the workspace.
Baselines: We compare HAND to the following baselines:
  • $\pi_{\text{base}}$: the base policy pre-trained on play data;
  • Lang Cond: $\pi_{\text{base}}$ with language-conditioning;
  • CLIP (Language): retrieves based on cosine similarity between language embedding of target task (rather than hand demo) and language embedding of the play data;
  • CLIP (Image): retrieves based on cosine similarity between language embedding of target task and image embedding of the play data;
  • Flow [12]: trains a VAE on pre-computed optical flows for $\mathcal{D}_{\text{play}}$ from GMFlow [44] and retrieves individual states-action pairs based on latent motion similarity; and
  • STRAP [13]: also uses S-DTW for sub-trajectory retrieval but uses S-DTW distance based solely on Euclidean distance between pre-trained DINO-v2 image embeddings.
STRAP and Flow assume access to expert robot demonstrations for both retrieval and fine-tuning. In our setting, we do not assume such demonstrations, therefore, unless otherwise noted, we adopt them without expert fine-tuning. While STRAP and Flow originally propose training policies from scratch, we instead apply LoRA fine-tuning—as with HAND —which we found to yield better performance for these baselines.
Policy Architecture: To ensure fair comparison, all methods use a three-layer action-chunking transformer (similar to ACT [36]) decoder policy where applicable. The input to the transformer policy is a sequence of image tokens corresponding to the external and over-shoulder camera views. Conditioned on the current image observation, the model predicts an action chunk corresponding to a second of execution.
Figure 3: WidowX Robot Arm Setup. We evaluate the scalability of HAND on 10 manipulation tasks on a WidowX robot arm in a kitchen setup [45].
Reach Green Block Push Button Close Microwave
Flow 7/25 0/25 0/25
STRAP 5/25 0/25 2/25
HAND(-VF) 9/25 13/25 9/25
HAND 15/25 18/25 11/25
TABLE I: Number of retrieved sub-trajectories performing demonstrated task. HAND retrieves more task-performing sub-trajectories than Flow and STRAP.
IV-B Experimental Evaluation
(Q1): HAND retrieves more task-relevant data. We analyze the quality of retrieved sub-trajectories between Flow, STRAP, and HAND. STRAP and HAND both use S-DTW-based trajectory retrieval, but STRAP relies purely on visual DINO-v2 embeddings for retrieval. We provide a single hand demonstration of three real robot tasks and retrieve the top $K=25$ matches from $\mathcal{D}_{\text{play}}$. Compared to STRAP and Flow, we observe in Table I that HAND retrieves more trajectories in which the robot performs the demonstrated task. STRAP relies exclusively on visual similarity, while Flow relies exclusively on motion similarity. Both methods struggle when there is a significant visual or motion gap between the target demonstrations (e.g., human hand videos) and the play dataset. In particular, for the Push Button task, STRAP is unable to retrieve any relevant trajectories in its top matches.
We also observe that visual filtering is necessary to retrieve trajectories where the target object is interacted with, as demonstrated by HAND(-VF), an ablation of HAND without visual filtering (Section III-C), having 30% worse retrieval performance than HAND in Table I.
Figure 4: Qualitative retrieval results on OOD scene. We visualize the top sub-trajectory match of Flow, STRAP, HAND without visual filtering (HAND(-VF)), and HAND on two OOD domain demonstrations recorded from an iPhone camera, showing approaching a K-Cup and putting it into the machine. Only HAND’s top match is relevant for both hand demonstrations.
Method $10^{\circ}$ Horiz. $20^{\circ}$ Horiz. $30^{\circ}$ Horiz. $10^{\circ}$ Vert $20^{\circ}$ Vert $30^{\circ}$ Vert
Flow 0 / 25 2 / 25 5 / 25 1 / 25 0 / 25 0 / 25
STRAP 1 / 25 10 / 25 13 / 25 12 / 25 11 / 25 11 / 25
HAND 21 / 25 18 / 25 19 / 25 16 / 25 13 / 25 14 / 25
TABLE II: Camera angle robustness results. # of relevant retrieved trajectories for the Put Lid on Pot task if we change the camera angle vertically and horizontally by $10^{\circ}$ increments. HAND retrieves $18\%$ more relevant trajectories compared to STRAP even in the extreme case of $30^{\circ}$ shift.
(Q2): HAND supports hand demonstrations from unseen environments and is robust to camera angle shifts. Because HAND retrieves based on relative hand motions, it is also effective with hand demonstrations from out-of-distribution (OOD) scenes. To illustrate, we collect hand demonstrations in a new environment using a handheld iPhone camera and a real coffee machine, while retrieving from robot play data recorded in a completely different scene with a toy coffee machine. In Figure 4, we show the lowest cost retrieved sub-trajectory of STRAP and Flow compared to HAND and a HAND ablation without the visual filtering step, HAND(-VF). Both of the retrieved trajectories for STRAP and Flow, along with the top trajectory for HAND(-VF) are irrelevant to the demonstrated task. For the first task, STRAP is able to retrieve the initial reaching motion toward the K-Cup but misses the crucial grasping segment, as it does not leverage motion for retrieval. Only HAND retrieves relevant robot trajectories for both hand demonstrations because it focuses on the motion demonstrated by the human hand after visual filtering.
Table II shows that HAND is also robust to shifts in camera angle for the Put Lid on Pot task, far more than Flow and STRAP. We measure the number of relevant retrieved trajectories of different methods after vertical and horizontal camera angle shifts of $10^{\circ}$ increments. In the most extreme setting of $30^{\circ}$ shift, HAND retrieves $18\%$ more relevant trajectories compared to STRAP. These camera angle shifts emulate head rotations on humanoid robots or camera movement on mobile manipulators, suggesting that HAND can even work in such settings where camera viewpoint change may occur.
Figure 5: Real-Robot Results. Task completion (including partial completion) out of 10 of $\pi_{\text{base}}$, STRAP, Flow, and HAND.
Method
Slide
Pot
Put Obj
in Pot
Put Lid on
Pot
Turn
Knob
Long
Horiz.
Lang Cond 2 0 0 0 0
CLIP (Language) 0 4 1 0 0
STRAP 0 0 2 0 0
HAND w/o Pre-training 1 1 1 2 0
HAND(-VF) 2 0 0 5 0
HAND(-CW) 2 5 3 5 0
HAND 5 6 4 6 3
HAND + Lang Cond 8 7 5 7 5
TABLE III: Long horizon Cook Carrot task results. We show success rates on each subtask and on the full task execution. Full Task Success is out of 10.
Figure 6: Fast Adaptation Study. We conduct a small-scale user study to demonstrate HAND’s ability to learn robot policies in real-time. From providing the hand demonstration (Left), to retrieval and fine-tuning a base policy (Middle), to evaluating the policy (Right), we show HAND can learn to put a carrot in the blender with 7.5/10 task completion in less than 4 minutes.
(Q3): HAND enables efficient policy learning in the real world. We evaluate four methods, including ours, across three standard tasks with ten trials each, for a total of 120 evaluations. Real-world experiments in Figure 5 on our three standard tasks demonstrate that fine-tuning with HAND improves success rates by +45% over the next best baseline, STRAP. In contrast, Flow fails to learn a policy that achieves reasonable success rates in any of the tasks. We also report the performance of $\pi_{\text{base}}$, trained on all of $\mathcal{D}_{\text{play}}$ and note that the pre-trained policy struggles to perform the tasks without any task-specific fine-tuning.
We next evaluate eight methods, including several ablations of HAND, across four base tasks and one long-horizon task constructed from these base tasks, with ten trials each for a total of 400 evaluations. Results on the more challenging long-horizon tasks (Table III) demonstrate that retrieval using hand demonstrations outperforms language-based retrieval (CLIP (Language)) by a factor of in success rate. Language-based retrieval suffers from the lack of spatial awareness, often retrieving trajectories that are semantically correct but spatially misaligned—similar to STRAP —which makes policy fine-tuning more difficult. In contrast, directly conditioning on language performs poorly compared to retrieval (Lang Cond), despite the fact that annotating sub-trajectories with language more than doubles data collection and annotation time.
Ablation Study: We observe that each component of HAND, namely pretraining, visual filtering (VF), and cost weighting (CW), are critical for task performance. Cost weighting helps bias the resulting policy towards behaviors that are most relevant to the downstream task, and reduces the effect of potentially noisy retrievals that may not directly aid in learning the target task. Without any of these components, the resulting policy is unable to learn the task. Only HAND successfully performs the entire long-horizon task reliably. We also show that given access to a language-annotated dataset, one could add language-conditioning on top of HAND to further improve the task performance (HAND + Lang Cond).
(Q4): HAND enables real-time, data-efficient policy learning of long-horizon tasks. Finally, we performed two small-scale user studies with IRB approval from our institution to demonstrate real-time learning. In the first study shown in Figure 6, a participant familiar with HAND iteratively demonstrated each part of a long-horizon Blend Carrot task and trained a HAND policy with over 70% success rate in under four (4) minutes from providing a single hand demonstration to deploying the fine-tuned policy. A video of a similar experiment can be found on our website.
Method User 1 (Minutes) User 2 (Minutes)
Hand Demonstrations (Min) $\downarrow$ 3 2
Robot Demonstrations (Min) $\downarrow$ 10 14
Hand Demonstrations (SR) $\uparrow$ 5/10 4/10
Robot Demonstrations (SR) $\uparrow$ 3/10 2.5/10
TABLE IV: Hand vs Robot Teleoperation. Time taken and success rates between hand and teleoperated demonstrations.
Hand vs Robot Demonstration Comparison: In the second study, two users with prior teleoperation experience—but not affiliated with this research—each collected a total of 20 demonstrations: 10 using hand teleoperation and 10 using robot teleoperation, to train the robot for the Put K-Cup in Coffee Machine task. We employ HAND retrieval for hand-collected demonstrations and STRAP retrieval for robot teleoperation demonstrations. For a direct comparison, we additionally fine-tune STRAP with the human-collected teleoperated demonstrations as per [13].
As reported in Table IV, teleoperated demonstrations required over $3\times$ more time to collect than hand demonstrations. Remarkably, with just a single hand demonstration per user, we fine-tuned a policy achieving over 40% task completion compared to STRAP which reaches only 25% using a single robot teleoperation demonstration. Interestingly, we observed that increasing the number of expert demonstrations used for STRAP degraded downstream performance likely due to lower quality retrieved trajectories. These results demonstrate that HAND enables fast adaptation to downstream tasks with as few as one (1) easy-to-provide hand demonstration.
V Conclusion and Limitations
We presented HAND, a simple and time-efficient framework for adapting robots to new tasks using easy-to-provide human hand demonstrations. We demonstrated that HAND enables real-time task adaptation with a single hand demonstration in under four minutes.
Extending to 3D paths for retrieval. While HAND uses 2D paths for retrieval, one future direction could extend HAND to estimate the hand trajectory in 3D using foundation depth prediction models. Another direction future work could consider is a mixture of features for improving retrieval for tasks that require more dexterous control, i.e., cloth folding or deformable object manipulation.
Severe Camera Viewpoint Changes. Future work could address issues from severe camera viewpoint shifts between the collected hand demonstrations and robot play data via the use of 3D information, multiple camera viewpoints, or scene re-rendering with virtual cameras [46].